OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4>

OpenStudy (ybarrap):

dot product: \[u*v=|u||v|\cos\] \[\cos \Theta =\frac{ u*v }{ |u||v|}\] \[\Theta = \cos^{-1} \frac{ u*v }{ |u||v| }\] \[<6,-1>*<7,-4>=<6*7,-1*-4>\] \[|u|=|<6,-1>|=\sqrt{6^{2}+-1^{2}}\] \[|v|=|<7,-4>|=\sqrt{7^{2}+-4^{2}}\] Plug-in and solve for inverse cosine (rounding your answer to 1/10 of degree)

OpenStudy (ybarrap):

Correction: \[<6,-1>*<7,-4>=6*7+(-1)*(-4)\]

OpenStudy (anonymous):

is it 20.3

OpenStudy (ybarrap):

yes